# Roman Garipov $^{*\,\dagger}$

### v2

@@ 5 Conclusion
In this work, we propose and evaluate a fully automated protocol for task-specific acceleration of speculative decoding. Our initial experiments suggest that a simple procedure can successfully determine which of the mismatching tokens in the LLM response affect the downstream quality for both mathematical reasoning and programming tasks. We hope that AutoJudge can facilitate the use of Judge Decoding across different task types, languages and modalities.
In future work, we aim to explore the performance of AutoJudge decoding across additional tasks and models and analyze how its outputs differ depending on the use case. Since our approach does not require human annotation, another research direction is to apply our algorithm to find important “tokens” for more complex openended tasks such as dialog or creative writing. While it is challenging to directly compare against the original Judge Decoding (see the end of Section 3), it would nonetheless be interesting to compare human and automated annotations in this setting. We also plan to explore how AutoJudge can be combined with more advanced speculative decoding algorithms, such as speculative decoding with tree-based drafts (Miao et al., 2023; Chen et al., 2024; Svirschevski et al., 2024) or learned drafting heads (Cai et al., 2024; Li et al., 2024b, 2025).
@@ Acknowledgements
To generalize our approach for stochastic generation, we take advantage of the well-know Gumbel-max trick [Gumbel, 1954]. To recap, if we add independent Gumbel-distributed random variables to each predicted logit and take the index of the maximum, the probability that a certain index will be chosen is equal to the softmax of the original logits.
In case of Algorithm 1, we use Gumbel-max trick to reparameterize stochastic sampling from the model with a deterministic sampling conditioned on a pre-generated random state $s\leftarrow\textsc{randBits}(N)$. Given a prompt $x$, a response prefix $y_{1:t}$ and model parameters $\theta$, we generate the next token as follows:
@@
where GumbelPRNG is a function that samples a pseudo-random variable from standard Gumbel distribution based on an input seed $s\oplus x\oplus y_{1:t}$. To recall, $\oplus$ denotes concatenation. This way, $y_{next}$ is distributed as $P(y_{next}|x\oplus y_{1:t},\theta)$, but it is deterministic when conditioned on the random state $s$. Hence, we sample a random state $s$ once at the beginning of Algorithm1, the entire procedure after that will also be conditionally deterministic (given $s$). We plan to test this in further versions of the paper.
@@
As we described earlier, Algorithm 1 is inherently sequential because it searches not for individual important tokens, but for important token combinations. In principle, it is tempting to consider a simpler algorithm that considers each token replacement in isolation and can run in parallel. However, when considering [target_model_gen_0, draft_token, target_model_gen_1] sequences only, a sufficiently strong target model might recover from even a low-quality token and still produce the correct answer. This results in a failure mode where all tokens are individually unimportant, but when all such tokens are jointly replaced with their draft versions, the model fails to produce the correct answer. In our preliminary experiments, when using Llama-3.1-70B-Instruct Touvron et al. [2023] as the target model and Llama-3.2-1B-Instruct as the draft model, fewer than 1% of the tokens were labeled as important with this simplified algorithm, whereas our main Algorithm 1 found substantially. One interesting guarantee of Algorithm 1 over its naïve counterpart is that, whenever draft and target models produce different (non-equivalent) answers to a given prompt, our algorithm will find at least one important token, whereas the naïve algorithm may find none.
@@ Appendix B Additional Details on Classifier Training
As we discussed earlier in Section 3.2, there are several important design choices that can affect the performance of an important token classifier in our setting. In this part of supplementary materials, we report the experiments that led us to use a linear classifier based on draft token embeddings encoded with both $\theta_{draft}$ and $\theta_{main}$. To that end, we compare the different classifier variants using the important token embeddings from the GSM8K [Cobbe et al., 2021] training subset (see Section 4.1).
To compare different classifier configurations, we further divide the GSM8K training set into classifier training (90%) and validation (10%) subsets. We perform this division at sample level, i.e. all labeled tokens from a given GSM8K sample are used either entirely for classifier training, or entirely for validation. We use the same training and validation subsets throughout this section.
For the first set of experiments (Figure 5), we compare regularizer coefficients for Logistic Regression (left). We also report different classifier types: Logistic Regression, a Random Forest with 128 trees and a multi-layer perceptron (MLP) with a single hidden layer consisting of 128 hidden units with ReLU activation. For consistency, we run all models using Scikit-Learn [Pedregosa et al., 2011] v1.4.2 with all other settings kept to their default values. For LogisticRegression we set max_iter parameter to $500$. For MLP, we perform early stopping on yet another 10% subset of the training set with built-in default MLPClassifier early stopping parameters. For RandomForest we use min_samples_leaf set to $0.001$. For this evaluation, all classifiers use draft and target model hidden states (concatenated) encoding the draft token, which is our main setup from Section 3.2.
@@ Appendix C Precision Matters for Speculative Decoding
This is important for AutoJudge since the token labeling Algorithm 1 runs sequential inference with $\theta_{target}$ and parallel inference on $\theta_{draft}$, whereas inference-time speculative decoding does it the other way around: sequential calculations of $\theta_{draft}$ during the draft generation phase, then parallel forward pass with $\theta_{target}$ during the verification phase. As a result, the classifier is trained on features that can be significantly different from what they would be during inference. In contrast, running in full precision (float32) does not have such problems.
@@ Appendix D Additional Evaluations
In Figure  8, we report accuracy and the number of accepted tokens for Llama-3.1-8B draft / Llama-3.1-70B target and Llama-3.1-8B draft / Llama-3.1-70B target model pairs in GSM8K 8-shot setup.
@@
To measure time efficiency of AutoJudge, we incorporate it into vLLM inference library. The integration is built upon vllm==0.8.5, torch==2.7.0 with CUDA 12.8 and transformers==4.51.3. We set window size $32$ and batch size $1$. The implementation allows to perform batched inference without modifications, but we choose to evaluate with batch size $1$ for accurate measurements. For efficiency the current implementation predicts important tokens using only hidden state of the target model, as we have shown that adding hidden states of the draft model does not substantially increase the accuracy. Evaluations for 1B/8B model pair and four for 8B/70B pair are run on NVIDIA A100-SXM4-80GB GPUs and on 8 NVIDIA H100-SXM5 80GB GPUs for 8B/405B in fp8 dtype format in tensor parallel setup.
Appendix F Additional Inference Benchmarks
In this section, we report additional runtime benchmarks that extend our evaluations from Section 4.3. We report our results for GSM8K with 1B/8B model pair in Table 2 and with 1B/70B (left) and 8B/70B (right) in Table 3. In both figures, we use the same experiment configuration as in Section 4.3.
For the 1B/8B configuration (Table 2), AutoJudge achieves throughput improvements of up to $1.83{\times}$ relative to standard speculative decoding. In the 1B/70B setup (Table 3, left), AutoJudge consistently surpasses speculative decoding in tokens per second while maintaining comparable accuracy. Similarly, in the 8B/70B setup (Table 3, right), AutoJudge achieves stable speedups, with relative improvements exceeding $1.3{\times}$ at optimal thresholds.
Table 2: Inference speed benchmarks on GSM8K (0-shot) for 1B draft / 8B target model with (left) and on GSM8K (8-shot) for 1B draft / 8B target model (right).
Llama 3.2 1B draft / 3.1 8B target
Threshold 0.06 0.09 0.12 0.16
Accuracy, % 83.1 82.0 80.2 77.4
Speed, tokens/s 116.3 124.9 134.0 154.9
Speculative Decoding:   84.2 tokens/s
Speedup(ours) 1.38 1.48 1.59 1.83
Llama 3.2 1B draft / 3.1 8B target
@@ Appendix F Additional Inference Benchmarks
Speed, tokens/s 86.4 94.2 96.7 99.1
Speculative Decoding:   65.2 tokens/s
Speedup(ours) 1.33 1.44 1.48 1.52
Table 3: Inference speed benchmarks on GSM8K 8-shot with vLLM implementation for 1B draft / 70B target models (left) and for 8B draft / 70B target models (right).
Llama 3.2 1B draft / 3.1 70B target
Threshold 0.0 0.01 0.03 0.05
Accuracy, % 94.7 95.1 95.3 94.6
Speed, tokens/s 47.9 76.1 77.8 76.6
Speculative Decoding:   47.9 tokens/s
Speedup(ours) 1.0 1.58 1.62 1.59
Llama 3.1 8B draft / 3.1 70B target
Threshold 0.05 0.11 0.18 0.28
Accuracy, % 94.8 93.4 92.9 90.4
Speed, tokens/s 64.4 75.9 75.6 81.6
@@
Speedup(ours) 1.59 1.87 1.86 2.01
@@ Appendix G Evaluation with Individually Tuned Window Sizes
Our main evaluations in Section 4.2 use a fixed window size of 32 tokens for consistency with our accepted token measurements. However, the optimal window size for every given setup depends on many factors: the exact draft/target model pair, GPU type and count, use of offloading and any implementation intricacies that may affect the ratio between drafting and verification phase latency.
In this section, we account for this by evaluating AutoJudge and speculative decoding with optimal window sizes for every model pair. We consider window sizes between 8 and 64 with the following values: [8,10,12,14,16,20,24,26,28,32,40,48,54,64][8,10,12,14,16,20,24,26,28,32,40,48,54,64] and report the best average tokens per second. Note that this evaluation protocol suffers from the natural variance in latency (e.g. due to varying individual clock rates) since we report the highest value measured.
The results for 1B/8B models pair are reported in Table  4. The results for 8B/70B pair are reported in Table 5.
Overall, AutoJudge decoding still outperforms standard speculative. Note however that the exact difference in terms of tokens per second will depend on the optimal window size, with setups that have higher target/draft forward-pass latency ratio (e.g. offloading) or lower lossless acceptance rates translating to higher performance advantage for AutoJudge.
Table 4: (left) Inference speed benchmarks on GSM8K 0-shot with vLLM implementation for 1B draft / 8B target models with tuned window size (Spec. Decoding = 8 tokens, Autojudge = 10 tokens) and (right) for 1B draft / 8B target models on GSM8K 8-shot with tuned window size (Spec. Decoding = 8 tokens, Autojudge = 8 tokens).
Llama 3.2 1B draft / 3.1 8B target
Threshold 0.06 0.09 0.12 0.16
Accuracy, % 83.1 82.0 80.2 77.4
Speed, tokens/s 164.2 168.4 171.1 178.3
@@
Speedup(ours) 1.11 1.14 1.15 1.20
Llama 3.2 1B draft / 3.1 8B target
@@
Speed, tokens/s 128.2 133.8 137.2 138.3
Autoregressive:   114.8 tokens/s
Speedup(ours) 1.12 1.17 1.19 1.20
Table 5: (left) Inference speed benchmarks on GSM8K 0-shot with vLLM implementation for 8B draft / 70B target models with tuned window size (Spec. Decoding = 8 tokens, Autojudge = 20 tokens) and (right) for 8B draft / 70B target models on GSM8K 8-shot with tuned window size (Spec. Decoding = 8 tokens, Autojudge = 16 tokens).
Llama 3.1 8B draft / 3.1 70B target
Threshold 0.009 0.015 0.031 0.145
Accuracy, % 92.3 92.3 91.9 89.9
Speed, tokens/s 82.7 81.4 94.1 99.1
Speculative Decoding:   70.9 tokens/s
Speedup(ours) 1.17 1.15 1.33 1.40
Llama 3.1 8B draft / 3.1 70B target
Threshold 0.03 0.11 0.18 0.28
Accuracy, % 95.4 93.4 92.9 90.4
Speed, tokens/s 64.9 71.2 75.6 74.3
@@ Appendix H Offloading
Llama 3.1 8B draft / 3.1 70B target
Threshold 0.03 0.05 0.11 0.18
Accuracy, % 95.4 94.8 93.4 92.9
Speed, tokens/s 0.7 0.7 0.9 1.1
Speculative Decoding:   0.62 tokens/s
Speedup(ours) 1.13 1.15 1.47 1.76

### v3

@@ 5 Conclusion
In this work, we propose and evaluate a fully automated protocol for task-specific acceleration of speculative decoding. Our experiments show that a simple procedure can automatically determine which of the mismatching tokens in the LLM response affect the downstream quality on a variety of tasks. Our runtime benchmarks demonstrate significant speedups on top of already tuned speculative decoding and EAGLE algorithms with minimal inference code modification. We hope that AutoJudge can facilitate the use of Judge Decoding across different task types, languages and modalities.
One promising direction for future research is to focus on open-ended problems such as instruction following or creative writing. While AutoJudge already demonstrates some speedups on open-ended tasks, it was primarily designed for technical problems such as math and programming. It would be interesting to investigate how it can be extended for longform generations with noisy and informal quality criteria. Another possible direction is to combine AutoJudge with additional speculative decoding algorithms, such as speculative decoding with tree-based drafts (Miao et al., 2023; Chen et al., 2024; Svirschevski et al., 2024) or learned drafting heads (Cai et al., 2024; Li et al., 2025).
@@ Acknowledgements
To generalize our approach for stochastic generation, we take advantage of the well-known Gumbel-max trick [Gumbel, 1954]. To recap, if we add independent Gumbel-distributed random variables to each predicted logit and take the index of the maximum, the probability that a certain index will be chosen is equal to the softmax of the original logits.
In case of Algorithm 1, we use the Gumbel-max trick to reparameterize stochastic sampling from the model as deterministic sampling conditioned on a predefined random state $s\leftarrow\textsc{randBits}(N)$. Given a prompt $x$, a response prefix $y_{1:t}$ and model parameters $\theta$, we generate the next token as follows:
@@
where GumbelPRNG is a function that samples a pseudo-random variable from standard Gumbel distribution based on an input seed $s\oplus x\oplus y_{1:t}$. To recall, $\oplus$ denotes concatenation. This way, $y_{next}$ is distributed as $P(y_{next}|x\oplus y_{1:t},\theta)$, but it is deterministic when conditioned on the random state $s$. Hence, we sample a random state $s$ once at the beginning of Algorithm1, the entire procedure after that will also be conditionally deterministic (given $s$).
@@
As we described earlier, Algorithm 1 is inherently sequential because it searches not for individual important tokens, but for important token combinations. In principle, it is tempting to consider a simpler algorithm that considers each token replacement in isolation and can run in parallel. However, when considering [target_model_gen_0, draft_token, target_model_gen_1] sequences only, a sufficiently strong target model might recover from even a low-quality token and still produce the correct answer. This results in a failure mode where all tokens are individually unimportant, but when all such tokens are jointly replaced with their draft versions, the model fails to produce the correct answer. In our preliminary experiments, when using Llama-3.1-70B-Instruct Touvron et al. [2023] as the target model and Llama-3.2-1B-Instruct as the draft model, fewer than 1% of the tokens were labeled as important with this simplified algorithm, whereas our main Algorithm 1 found substantially more. One interesting guarantee of Algorithm 1 over its naïve counterpart is that whenever draft and target models produce different (non-equivalent) answers to a given prompt, our algorithm will find at least one important token, whereas the naïve algorithm may find none.
@@ Appendix B Additional Details on Classifier Training
As we discussed earlier in Section 3.2, there are several important design choices that can affect the performance of an important token classifier in our setting. Here, we report the experiments that led us to use a linear classifier based on draft token embeddings encoded with both $\theta_{draft}$ and $\theta_{main}$. To that end, we compare the different classifier variants using the important token embeddings from the GSM8K [Cobbe et al., 2021] training subset (see Section 4.1).
To compare different classifier configurations, we further divide the GSM8K training set into classifier training (90%) and validation (10%) subsets. We perform this division at the sample level, i.e., all labeled tokens from a given GSM8K sample are used either entirely for classifier training, or entirely for validation. We use the same training and validation subsets throughout this section.
For the first set of experiments (Figure 6), we compare regularizer coefficients for Logistic Regression (left). We also report different classifier types: Logistic Regression, a Random Forest with 128 trees and a multi-layer perceptron (MLP) with a single hidden layer consisting of 128 hidden units with ReLU activation. For consistency, we run all models using scikit-learn [Pedregosa et al., 2011] v1.4.2 with all other settings kept to their default values. For LogisticRegression, we set max_iter parameter to $500$. For MLP, we perform early stopping on yet another 10% subset of the training set with the built-in default MLPClassifier early stopping parameters. For RandomForest, we use min_samples_leaf set to $0.001$. For this evaluation, all classifiers use draft and target model hidden states (concatenated) encoding the draft token, which is our main setup from Section 3.2.
@@ Appendix C Precision Matters for Speculative Decoding
This is important for AutoJudge, since Algorithm 1 runs sequential inference with $\theta_{target}$ and parallel inference on $\theta_{draft}$, whereas inference-time speculative decoding does it the other way around: sequential calculations of $\theta_{draft}$ during the draft generation phase, then parallel forward pass with $\theta_{target}$ during the verification phase. As a result, the classifier is trained on features that can be significantly different from what they would be during inference. In contrast, running in full precision (float32) does not have such problems.
@@ Appendix D Additional Evaluations
In Figure  9, we report accuracy and the number of accepted tokens for Llama-3.1-8B draft / Llama-3.1-70B target and Llama-3.1-8B draft / Llama-3.1-70B target model pairs in GSM8K 8-shot setup.
@@
To measure time efficiency of AutoJudge, we incorporate it into vLLM inference library. The integration is built upon vllm==0.8.5, torch==2.7.0 with CUDA 12.8 and transformers==4.51.3. We use a window size of $32$ and a batch size of $1$. The implementation allows to perform batched inference without modifications, but we choose to evaluate with batch size $1$ for accurate measurements. For efficiency, the current implementation predicts important tokens using only the hidden state of the target model, as we have shown that adding hidden states of the draft model does not substantially increase the accuracy. We run evaluations for the 1B/8B model pair and for the 8B/70B pair 1 and 4 NVIDIA A100-SXM4-80GB GPUs respectively and on 8 NVIDIA H100-SXM5 80GB GPUs for 8B/405B in FP8 precision in the tensor parallel setup.
Appendix F Additional Inference Benchmarks: Equal Window Size
In this section, we report additional runtime benchmarks that extend our evaluations from Section 4.3. For these experiments, we use a fixed window size of 32 for both vanilla speculative decoding and AutoJudge (as opposed to individually tuned window size in Section 4.3).
We report our results for GSM8K with 1B/8B model pair in Table 3, 8B/70B in Table 4 and 1B/70B in Table 5. All three model pairs show significant speedups relative to standard speculative decoding. Notably, our 8B/70B setup has up to 2.5x with about 2.5% loss in accuracy.
Table 3: Inference speed benchmarks on GSM8K for 1B draft / 8B target model on: (left) GSM8K 0-shot evaluation and (right) GSM8K 8-shot evaluation.
Llama 3.2 1B draft / 3.1 8B target (0-shot)
Threshold 0.04 0.09 0.12 0.16
Accuracy, % 83.9 82.0 80.2 77.4
Speed, tokens/s 89.1 106.8 121.0 146.2
Speculative Decoding:   84.5 tokens/s
Speedup(ours) 1.05 1.26 1.43 1.72
Llama 3.2 1B draft / 3.1 8B target (8-shot)
@@ Appendix F Additional Inference Benchmarks: Equal Window Size
Speed, tokens/s 66.5 74.3 76.3 83.3
Speculative Decoding:   62.4 tokens/s
Speedup(ours) 1.07 1.19 1.22 1.33
Table 4: Inference speed benchmarks on GSM8K for 8B draft / 70B target model on: (left) GSM8K 0-shot evaluation and (right) GSM8K 8-shot evaluation.
Llama 3.1 8B draft / 3.1 70B target (0-shot)
Threshold 0.005 0.031 0.145 0.230
Accuracy, % 92.0 91.9 89.9 88.0
Speed, tokens/s 41.8 80.6 107.4 109.5
Speculative Decoding:   40.6 tokens/s
Speedup(ours) 1.03 1.98 2.64 2.70
Llama 3.1 8B draft / 3.1 70B target (8-shot)
Threshold 0.03 0.05 0.18 0.28
Accuracy: Accuracy, % 95.4 94.8 92.9 90.4
Speed, tokens/s 64.4 71.1 79.3 86.2
@@
Speedup(ours) 1.58 1.74 1.94 2.11
Table 5: Inference speed benchmarks on GSM8K 8-shot with vLLM implementation for 1B draft / 70B target models.
Llama 3.2 1B draft / 3.1 70B target (8-shot)
Threshold 0.01 0.03 0.05 0.11
Accuracy, % 95.1 95.3 94.6 92.3
Speed, tokens/s 50.1 65.6 75.5 79.9
Speculative Decoding:   45.7 tokens/s
Speedup(ours) 1.10 1.44 1.65 1.75
@@ Appendix G Evaluation with Individually Tuned Window Sizes
As we discussed in Section 4.3, we tune the speculation window size individually for AutoJudge and vanilla speculative decoding to provide a more competitive baseline. This is because traditional speculative decoding accepts, on average, less tokens and does not benefit from having a larger draft window size. Thus, the two algorithms often work best with different draft sizes.
In this section, we account for this by evaluating AutoJudge and speculative decoding with optimal window sizes for every model pair. We consider window sizes between 6 and 64 with the following values: [6,8,10,12,14,16,20,24,26,28,32,40,48,54,64][6,8,10,12,14,16,20,24,26,28,32,40,48,54,64] and choose the best window size in terms of tokens per second. Note that this evaluation protocol suffers from the natural variance in latency (e.g. due to varying individual clock rates) since we report the highest value measured.
The results for 1B/8B models pair are reported in Table  6. The results for 8B/70B pair are reported in Table 7. Overall, AutoJudge decoding still significantly outperforms standard speculative decoding.
Table 6: (left) Inference speed benchmarks on GSM8K 0-shot with vLLM implementation for 1B draft / 8B target models with tuned window size (Spec. Decoding = 8 tokens, AutoJudge = 10 tokens) and (right) for 1B draft / 8B target models on GSM8K 8-shot with tuned window size (Spec. Decoding = 6 tokens, Autojudge = 10 tokens).
Llama 3.2 1B draft / 3.1 8B target (0-shot)
Threshold 0.06 0.12 0.15 0.16
Accuracy, % 83.1 80.2 79.8 77.4
Speed, tokens/s 149.2 169.2 171.2 173.9
@@
Speedup(ours) 1.01 1.14 1.15 1.17
Llama 3.2 1B draft / 3.1 8B target (8-shot)
@@
Speed, tokens/s 114.4 125.0 132.1 139.3
Speculative Decoding:   116.8 tokens/s
Speedup(ours) 0.98 1.07 1.13 1.19
Table 7: (left) Inference speed benchmarks on GSM8K 0-shot with vLLM implementation for 8B draft / 70B target models with tuned window size (Spec. Decoding = 8 tokens, Autojudge = 32 tokens) and (right) for 8B draft / 70B target models on GSM8K 8-shot with tuned window size (Spec. Decoding = 8 tokens, Autojudge = 16 tokens).
Llama 3.1 8B draft / 3.1 70B target (0-shot)
Threshold 0.005 0.031 0.145 0.230
Accuracy, % 92.0 91.9 89.9 88.0
Speed, tokens/s 72.3 80.6 107.4 109.6
Speculative Decoding:   72.3 tokens/s
Speedup(ours) 1.0 1.11 1.49 1.52
Llama 3.1 8B draft / 3.1 70B target (8-shot)
Threshold 0.03 0.05 0.18 0.28
Accuracy, % 95.4 94.8 92.9 90.4
Speed, tokens/s 69.0 73.1 83.6 84.5
@@ Appendix H Offloading
Llama 3.1 8B draft / 3.1 70B target (8-shot)
Threshold 0.03 0.05 0.11 0.28
Accuracy, % 95.4 94.8 93.4 90.4
peed, tokens/s 1.4 1.6 1.9 2.4
Speculative Decoding:   1.19 tokens/s
Speedup(ours) 1.20 1.31 1.59 1.98